Find the general solution of the differential equation nts y – 7y +12y = 2te3+ +1 - t y(t) = 20e3+ + boe4 – 2te3 – 12e31 -
Answer #1

y74tiny = 2te3t Its homogeneous equation -490 Its auxiliary equation is, .4m-8m + 12 = 0 (-4)-3(-4)=0 (-4) (-3)=0 are reay, act) at 109) et (atestttt) Slate**+-+) + <t) ++ Stetat & Jetta - Steutat Sletet arkt 2=2[-6.J-45 = -ztet_zé - 1 2 3 4 + t